# 2D Entanglement from a Twisted Motzkin Spaghetto

### v2

@@ IV Antiferromagnetic order
The expectation value of the spin in a 1D chain has been computed for $q=1$ from combinatorial methods [30, 31], field theoretic approach in the scaling limit [29], and the transfer matrix method [27]. All three approaches lead to the same scaling behavior as $r^{-1/2}$, where $r$ is the distance from the closer boundary, with up spin on the first half, and down spin on the second. When twisted and embedded on the 2D lattice, an induced antiferromagnetic order appears with modulated amplitude that decays from the boundary to the center, as shown in 2 (b).
As was shown in Ref. [27], due to the opposite boundary conditions, the $q>1$ ordered phase forms a domain wall between the two regions with positive and negative values of the order parameter. The correlation length in this case is reflected in the domain wall thickness $\xi$ by
@@
which can be derived from the Ginzburg-Landau theory [27]. The length scale depends on the deformation parameter as $\xi\propto 1/\log q$, which was obtained from numerical results by a matrix product state calculation [27]. Fig. 2 (a) depicts the twisting induced antiferromagnetic order in the direction orthogonal to 1D chain within each quadrant of the 2D lattice for $\xi=2N$.
V TN
@@
The TN representation of the Motzkin and Fredkin GS defined in [35] actually gives an inverted step pyramid shaped network with the physical external legs at the bottom of each column of tensors. In this section, we define a new TN representation of the Fredkin GS with an upright pyramid shaped geometry, which consist of slightly sparser tensors. This TN is obtained using the tiling approach used in Ref. [35]. In this method the correspondence between spin configurations and walks is used to construct the TN representation. This is achieved through an intermediate step where a valid tiling of a set of tiles is associated to each walk present in the GS. In this context, a valid tiling is a tiling where each edge of each tile matches its adjacent tiles. By carefully choosing an appropriate set of tiles, one can assure a one-to-one correspondence between valid tilings and spin configurations present in the GS superposition. For spin configurations present in the GS of the Fredkin model, which are described by Dyck walks, this one-to-one correspondence can be assured by the set of tiles $B_{i}(\bm{c})$ seen in panel (a) of figure 4. In panel (b) and (c) we see that the valid tilings contain arrowed paths indicating the color correlation between spins in the given Dyck walk. A colored spin up (resp. down) at $i$ is represented by a colored arrowed path entering (resp. leaving) the tiling at the floor edge of the bottom tile in the tower of tiles at $i$. Clearly, each valid tiling corresponds to a unique way of pairing the degrees of freedom. This set of tiles enables exactly one valid tiling for any given Dyck walk and no valid tiling for non-Dyck walks. To see this, note that by following the highest arrowed path within each tiling, the corresponding Dyck walk is traced out, see figure 4 (b) and (c). It is clear that a valid tiling for a given Dyck walk where this is not the case, is not possible. It is also clear that no valid tiling can be constructed for a non-Dyck walk, as that would necessarily violate arrow continuity at the step where the walk becomes negative (or at the endpoints if the walk starts or terminates at a non-zero height). This establishes the one-to-one correspondence between valid tilings and the spin configurations in the ground state.
The tensor building up the TN representation is then defined by the set of tiles in the following way: The number of indices of the tensor is determined by the shape of the tiles and each tile represents an index configuration of the tensor giving a nonzero value. Contractions between two tensors in the TN, then correspond to summing over tile configurations matching at that edge. The tiles, valid tilings and the resulting TN, can be seen in figure 4.
@@
Figure 4: (a) The five different tiles $B_{i}(\bm{c})$ for the new Fredkin tensor network. The tiles are also defined as rank-4 tensors in terms of Kronecker deltas, where the indices $\bm{k}_{i}$ are defined as for $B(q)$ in panel (d). We have $\bm{c}=(1,0)$ for red arrow, $\bm{c}=(0,1)$ for blue arrow. No arrow corresponds to $\bm{0}$. (b)-(c) Valid tilings corresponding to the maximal and minimal height Dyck walk for the $L=4$ system. (d) TN representation of the ground state of the single Fredkin chain for $L=4$. The constituent tensors are defined in the upper panel. $l$ denotes the different levels of the holographic TN and the white circle projects out the index value $\bm{0}$ corresponding to no arrow in the tilings.
@@
$\begin{split}B(q)=\sum_{\bm{c}_{i}=(1,0),(0,1)}&[\sqrt{q}B_{1}(\bm{c}_{1})+qB_{2}(\bm{c}_{1},\bm{c}_{2})\\ +&qB_{3}(\bm{c}_{1},\bm{c}_{2}),+\sqrt{q}B_{4}(\bm{c}_{1})+B_{5}(\bm{c}_{1})].\end{split}$ (5)
The indices $\bm{k}_{1},\cdots,\bm{k}_{4}$ of $A(q)$ are defined in panel (d) of figure 4 and have been suppressed in equation 5. We use a vector notation for the indices to keep track of the color of the arrows, which is a notation conveniently extended to the tensors in the TN for the 2D systems [29]. It is clear that one can simply replace the different vectors with a unique number, to arrive at the more familiar way of denoting index configurations by numbers. Note also that we use a minus sign to indicate arrows leaving (resp. coming in) at the $\bm{k}_{1}$ (resp. $\bm{k}_{3}$) index. In equation 5, we consider two colors, namely red, $\bm{c}=(1,0)$, and blue, $\bm{c}=(0,1)$. Note that each of the tiles $B_{i}(\bm{c})$ can be defined as a rank-4 tensor expressed as a product of Kronecker deltas, as seen in panel (a) in figure 4. The shape of the resulting TN naturally becomes that of the valid tilings, giving an upright step pyramid shape as seen in panel (d) in figure 4. Note that the boundary tensors $\delta_{\bm{k},\bm{0}}$ ensures that “no arrows flow out” of the TN, except at the bottom edges. Due to the tile without arrows, we must project out the index value corresponding to an empty edge at the bottom tile. We must therefore include projectors in the TN, as seen in figure 4 (d). This is similar to the rainbow TN presented with projectors in Ref. [35].
@@
VI Conclusion
In this article, a frustration free 2D lattice Hamiltonian with exactly solvable GS that has volume scaling of EE is constructed by twisting the highly entangled Motzkin chain. The interaction is significantly simpler than the previous and only model of such highly entangled 2D systems realized by coupling arrays of highly entangled chains. Similar constructions of 2D entanglement from twisting 1D chains can be applied to the highly entangled XX chain with inhomogeneous coupling [8, 9, 10]. At the moment, it is not clear if similar construction works for 3D systems, whereas the coupled chains construction in principle works in any dimension via tiling. These alternative examples of area-law violating GSs provide a better understanding of the interplay between EE and spectral gap scaling, or the tradeoff between quantum computing resource and robustness against noise. The ultimate goal in this direction is to generalize the area law to gapless ground states to get a relation between the rate of decay in spectral gap and growth of EE with the system size.
Acknowledgements.

### v3

@@ 4 Antiferromagnetic order in the radial direction
The expectation value of the spin in a 1D chain has been computed for $q=1$ from combinatorial methods [31, 32], field theoretic approach in the scaling limit [30], and the transfer matrix method [28]. All three approaches lead to the same scaling behavior as $r^{-1/2}$, where $r$ is the distance from the closer boundary, with up spin on the first half, and down spin on the second. When twisted and embedded on the 2D lattice, an induced antiferromagnetic order appears with modulated amplitude that decays from the boundary to the center, as shown in 3 (b).
As was shown in Ref. [28], due to the opposite boundary conditions, the $q>1$ ordered phase forms a domain wall between the two regions with positive and negative values of the order parameter. The correlation length in this case is reflected in the domain wall thickness by
@@
[…]
5 TN representation of the GS
@@
The TN representation of the Motzkin and Fredkin GS defined in [36] actually gives an inverted step pyramid shaped network with the physical external legs at the bottom of each column of tensors. In this section, we define a new TN representation of the Fredkin GS with an upright pyramid shaped geometry, which consist of slightly sparser tensors. This TN is obtained using the tiling approach used in Ref. [36]. In this method the correspondence between spin configurations and walks is used to construct the TN representation. This is achieved through an intermediate step where a valid tiling of a set of tiles is associated to each walk present in the GS. In this context, a valid tiling is a tiling where each edge of each tile matches its adjacent tiles. By carefully choosing an appropriate set of tiles, one can assure a one-to-one correspondence between valid tilings and spin configurations present in the GS superposition. For spin configurations present in the GS of the Fredkin model, which are described by Dyck walks, this one-to-one correspondence can be assured by the set of tiles $B_{i}(\bm{c})$ seen in panel (a) of figure˜5. In panel (b) and (c) we see that the valid tilings contain arrowed paths indicating the color correlation between spins in the given Dyck walk. A colored spin up (resp. down) at $i$ is represented by a colored arrowed path entering (resp. leaving) the tiling at the floor edge of the bottom tile in the tower of tiles at $i$. Clearly, each valid tiling corresponds to a unique way of pairing the degrees of freedom. This set of tiles enables exactly one valid tiling for any given Dyck walk and no valid tiling for non-Dyck walks. To see this, note that by following the highest arrowed path within each tiling, the corresponding Dyck walk is traced out, see figure˜5 (b) and (c). It is clear that a valid tiling for a given Dyck walk where this is not the case, is not possible. It is also clear that no valid tiling can be constructed for a non-Dyck walk, as that would necessarily violate arrow continuity at the step where the walk becomes negative (or at the endpoints if the walk starts or terminates at a non-zero height). This establishes the one-to-one correspondence between valid tilings and the spin configurations in the ground state.
The tensor building up the TN representation is then defined by the set of tiles in the following way: The number of indices of the tensor is determined by the shape of the tiles and each tile represents an index configuration of the tensor giving a nonzero value. Contractions between two tensors in the TN, then correspond to summing over tile configurations matching at that edge. The tiles, valid tilings and the resulting TN, can be seen in figure˜5.
@@
Figure 5: (a) The five different tiles $B_{i}(\bm{c})$ for the new Fredkin tensor network. The tiles are also defined as rank-4 tensors in terms of Kronecker deltas, where the indices $\bm{k}_{i}$ are defined as for $B(q)$ in panel (d). We have $\bm{c}=(1,0)$ for red arrow, $\bm{c}=(0,1)$ for blue arrow. No arrow corresponds to $\bm{0}$. (b)-(c) Valid tilings corresponding to the maximal and minimal height Dyck walk for the $L=4$ system. (d) TN representation of the ground state of the single Fredkin chain for $L=4$. $l$ denotes the different levels of the holographic TN. (e) The constituent tensors of the TN, where the white circle projects out the index value $\bm{0}$ corresponding to no arrow in the tilings.
@@
$\begin{split}B(q)=\sum_{\bm{c}_{i}=(1,0),(0,1)}&[\sqrt{q}B_{1}(\bm{c}_{1})+qB_{2}(\bm{c}_{1},\bm{c}_{2})+qB_{3}(\bm{c}_{1},\bm{c}_{2}),+\sqrt{q}B_{4}(\bm{c}_{1})+B_{5}(\bm{c}_{1})].\end{split}$ (11)
The indices $\bm{k}_{1},\cdots,\bm{k}_{4}$ of $A(q)$ are defined in panel (d) of figure˜5 and have been suppressed in equation˜11. We use a vector notation for the indices to keep track of the color of the arrows, which is a notation conveniently extended to the tensors in the TN for the 2D systems [30]. It is clear that one can simply replace the different vectors with a unique number, to arrive at the more familiar way of denoting index configurations by numbers. Note also that we use a minus sign to indicate arrows leaving (resp. coming in) at the $\bm{k}_{1}$ (resp. $\bm{k}_{3}$) index. In equation˜11, we consider two colors, namely red, $\bm{c}=(1,0)$, and blue, $\bm{c}=(0,1)$. Note that each of the tiles $B_{i}(\bm{c})$ can be defined as a rank-4 tensor expressed as a product of Kronecker deltas, as seen in panel (a) in figure˜5. The shape of the resulting TN naturally becomes that of the valid tilings, giving an upright step pyramid shape as seen in panel (d) in figure˜5. Note that the boundary tensors k,0 ensures that “no arrows flow out” of the TN, except at the bottom edges. Due to the tile without arrows, we must project out the index value corresponding to an empty edge at the bottom tile. We must therefore include projectors in the TN, as seen in figure˜5 (d). This is similar to the rainbow TN presented with projectors in Ref. [36].
@@
6 Conclusion
In this article, a frustration free 2D lattice Hamiltonian with exactly solvable GS that has volume scaling of EE is constructed by turning on the spatially anisotropic interaction following a specific path along which the degrees of freedom form a highly entangled Motzkin chain. The interaction is significantly simpler than the previous models of such highly entangled 2D systems realized by coupling arrays of highly entangled chains. Similar constructions of 2D entanglement from twisting 1D chains can be applied to the highly entangled XX chain with inhomogeneous coupling [8, 9, 10]. The volume scaling of EE in 2D is possible since the entangling color degree of freedom tend to form Bell pairs with their point reflection partner about the center of the system. Hence, at the moment, it is not clear if similar nontrivial construction works for 3D systems, for which an inversion symmetric path traversing all lattice points in the 3D space has to be specified. On the other hand, the coupled chains construction in principle works in any dimension via tiling. These alternative examples of area-law violating GSs provide a better understanding of the interplay between EE and spectral gap scaling, or the tradeoff between quantum computing resource and robustness against noise. The ultimate goal in this direction is to generalize the area law to gapless ground states to get a relation between the rate of decay in spectral gap and growth of EE with the system size.
Acknowledgements
@@
Appendix A Typical height change in a segment of Motzkin path
In this appendix, we determine the typical height change in a segment of Motzkin path by finding the most probable height change using combinatorics. The number of Motzkin path segments that start at height $h_{i}$ and end at height $h_{f}$ over a horizontal distance of $x$ can be enumerated using André’s reflection method, see figure˜6. Such paths meeting the Motzkin condition of non-negative height everywhere is equal to the total number of random walks subtracted by those that does reach height 1-1 somewhere. The latter are in one-to-one mapping with paths that are mirror images about the horizontal line h=1h=-1 after the first time the path reaches height 1-1. These paths are in turn enumerated by the random walks from height $h_{i}$ to height $h_{f}-2$.
Figure 6: The number of segments of Motzkin paths that start at height $h_{i}$ and end at height $h_{f}$ over $x$ steps without having negative heights in the middle can be enumerated with André’s reflection method.
For simplicity, we first count such Dyck paths, which are Motzkin paths without the option of flat moves, as
[…]
where the limit $x\gg h_{i},h_{f}\gg 1$ has been used. The most probably height change $h_{f}^{*}-h_{i}$ is determined by $p_{\mathrm{D}}^{\prime}(h_{f}^{*},h_{i},x)=0$, which is solved from
$e^{\frac{4h_{i}h_{f}^{*}}{x}}=\frac{h_{f}^{*}+h_{i}}{h_{f}^{*}-h_{i}}.$ (A.2)
Motzkin paths starting and ending at finite height can be counted analogously, resulting in
$\sqrt{p_{\mathrm{M}}(h_{f},h_{i},x)}\propto e^{-\frac{3(h_{f}-h_{i})^{2}}{2x}}-e^{-\frac{3(h_{f}+h_{i})^{2}}{2x}},$ (A.3)
and the typical height change determined by
$e^{\frac{6h_{i}h_{f}^{*}}{x}}=\frac{h_{f}^{*}+h_{i}}{h_{f}^{*}-h_{i}}.$ (A.4)
The leading contribution to EE is proportional to the total net height change between the two subsystems
$S\sim\sum_{i=0}^{\frac{L}{2}-1}|\langle h_{2i+1}-h_{2i}\rangle|\propto\int\mathcal{D}[h]\left[\sum_{k=1}^{\frac{L}{2}}(-1)^{k}h_{k-1}+\sum_{k=\frac{L}{2}}^{L-1}(-1)^{k}h_{k}\right],$ (A.5)
where $h_{i}$ are the height variable at the $L$ endpoints of the segments resulting from the bipartition on the 2D lattice. In the second form of the equation the contribution from the segments lying in the first half of the chain and contributions from segments lying in the second half of the chain have been split into two sums, using that in the first half of the chain the height increases, while in the second half of the chain it decreases. Throughout the paper, $L$ assumed an integer multiple of 4 without loss of generality. The functional integral over the height configuration $h(x)$ is done by integrating over the height variable at discrete locations weighted by corresponding probabilities
$\mathcal{D}[h]=p(h_{1})\prod_{i=1}^{\frac{L}{2}-1}dh_{i}p(h_{i+1},h_{i},|4i-2L|)\prod_{j=\frac{L}{2}}^{L-2}dh_{j}p(h_{j+1},h_{j},|4(j+1)-2L|)p(h_{L-1}),$ (A.6)
with the probability distribution of boundary segments are given by [11, 14]
[…]